AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
nToF
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri Apr 4 22:28:14 2025
Author
*
:
Subject
*
:
<p>E front 2x16-way to 1x34-way IDC cable from the Junction Box to the IDC-Lemo 00 adaptor was found to be inserted upside-down - the 16-way IDC connectors did not have polarisation keys - meaning the E detector p+n junction strip signals collected with 26Al for the first night are not working. Signals from the dE p+n junction strips and E n+n Ohmic strips will be OK. The mistake occurred at the conclusion of the noise tests yesterday - see https://elog.ph.ed.ac.uk/nToF/52</p> <p>The 2x16-way to 1x34-way IDC cable without polarisation keys has now been replaced.</p> <p>Run 216417 and onwards have correct cabling. For runs before that, E back and dE are still reliable.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Attachment 1:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown